# Attribute First, then Generate: Locally-attributable Grounded Text Generation

### v1

@@ C.1 Content Selection
In the ICL method, we present the model with the source texts and guide it to copy specific segments from those texts (see Fig. B.1 for a prompt example in the MDS setting). We use a designated “<SPAN_DELIM>” markup as a delimiter between such consecutive spans, for better parsing. Within the MDS framework, the model is tasked with identifying salient excerpts, whereas in the context of LFQA, the model is also provided with a query, and is asked to find the parts of the texts that are relevant to answering this question. Furthermore, following our analyses of each setting’s development sets, we restrict the model to an average and maximum number of words to choose (an average of 200 and 100 words and a maximum of 900 and 200 words for MDS and LFQA, respectively). Lastly, to manage source texts that exceed the model’s input size limit, we truncate the fewest necessary sentences from those texts to ensure they fit within the specified limit. [^17] We use spaCy Honnibal and Montani (2017) to separate documents into sentences.
After performing hyperparameter tuning, in the MDS we set $T=0.1$ and the number of in-context examples to 2, while in LFQA we set $T=0.3$ and the number of examples to 4.
@@
An example is provided in Fig. B.5. The input to the model is the set of all documents, concatenated with a special markup “<doc-sep>”, which was also used in pre-training in Primera. To fit the input in a context length of 4096, we trim each document to the length of 4096 divided by the number of documents in each example. The output is a set of highlights separated with a new special markup “<highlights-separator>”. We run hyperparameter-tuning on each setting based on the development set. The hyperparameters used for MDS are warmup steps of 229, weight decay of 0.5, and a learning rate of 4e-4. Conversely, the hyperparameters used for LFQA are warmup steps of 234, weight decay of 0, and a learning rate of 1e-4. We chose the best model checkpoint by calculating Intersection-over-Union (IoU) between the gold highlights spans and the predicted highlights spans.
@@ C.2 Generation Planning
In the ICL setting, we present the model with the source text where we insert the mark-ups “<highlight_start>” and “<highlight_end>” before and after each pre-selected span, (see Fig. B.2 for the MDS setting). We also enumerate the highlights across all documents, and for the LFQA setting, we also add the query. The model is directed to cluster highlighted segments that are likely to co-occur in a sentence in the same order it would then merge them into an output. We discovered that providing the model with task-specific motivation enhances its performance; for instance, in MDS, the objective is to fuse highlights into a summary, whereas in LFQA, it aims to formulate a response to the query. To facilitate the parsing of the model’s response, we encourage it to generate its response as a list of single-value dictionaries, with “cluster” as the key and a list of indices corresponding to the highlights within the current cluster as the value. For hyperparameter tuning, we divide the gold “highlights” into pairs. Pairs that correspond to the same target output sentence are categorized as positive, whereas others are considered negative. Subsequently, we evaluate the F1 score to assess the alignment between the predicted and gold clusters. This involves calculating the True Positives, False Negatives, and False Positives for the highlights pairs, based on the predicted clusters. Following our analyses, we opt for 2 examples and $T=0.5$ in the MDS setting, and 1 example and $T=0.1$ in the LFQA setting.
@@
An example is provided in Fig. B.6. There are two variants for generation planning: generative clustering and joint. In the generative clustering approach, the input to the model is the set of all highlights, concatenated with a special markup “<highlight-separator>”, which is the same as the output from the content selection model. The output is then a set of clusters, where each cluster is separated by a special markup “<cluster-separator>”, and inside each cluster we have a set of highlights in the same format as the input. We run hyperparameter-tuning on each setting based on the development set. The hyperparameters used for MDS are warmup steps 0, weight decay 0, and learning rate 5e-5. The hyperparameters used for LFQA are warmup steps 298, weight decay 0, and learning rate 2e-4. At inference time, we used a beam size of 1. We use the same logits processor described in Section C.1 to enforce the model to generate valid highlights.
@@ C.3 Attributed Fused Output
For the ICL variant, we iteratively show the model only the current cluster’s highlights, in the same manner, we do so in the generation planning ICL experiments (see Section C.2). Additionally, we present the model with what it generated to that point and guide it to generate the next sentence in the paragraph, which should cover all and only the highlighted spans. We use 3 examples and $T=0.3$ for the MDS setting, and 2 examples with $T=0.1$ for the LFQA setting.
For the Chain-of-Thought approach, which merges generation planning with the fusion steps, we present the model with the “highlighted” source texts and ask it to fuse those highlights into a coherent output. Additionally, we guide the model, both in the instructions and through in-context demonstrations to iteratively cluster highlights and fuse them into a sentence, before generating the final output (see Fig. B.4). We find that 3 in-context examples and $T=0.1$ work best for MDS, while for LFQA it is 1 example and $T=0.3$ that yield the best results on the development set.
@@
An example is provided in Fig. B.7. The input to the attributable fusion model is a single cluster of highlights, where highlights are separated by a special markup. The output is then a single sentence of the response. This model is called iteratively, and at each iteration, we include the prefix of the already generated response. The output also includes citations for every sentence in the generated response based on the highlights used to generate the particular sentence. The hyperparameters for the fusion model are for MDS are warmup steps 54, weight decay 0.2, and learning rate 1e-4. The hyperparameters used for LFQA are warmup steps 175, weight decay 0.5, and learning rate 2e-5.
@@ C.4 End-to-end Generation
We include as a baseline end-to-end generation without decomposing the process into multiple steps.
@@
For the ICL setting, we instruct the model to perform the corresponding setting’s task, namely to summarize the given documents for MDS, and to answer the query based on the supplied document for LFQA. In addition to the varying number of in-context examples, we also evaluate the model in a zero-shot setting. Our analyses show that for MDS, 1 example and $T=0.5$ give the best performance, while for LFQA 3 examples and $T=0.7$ are optimal.
@@ C.6 Constrained Decoding
Both the content selection component and the planning component are designed to exactly copy text spans from source documents. To guide the generation to precisely follow the lexical structure of the input’s text, we employ a constrained decoding strategy for the fine-tuned models. This strategy relies on a logits processor [^18] https://huggingface.co/docs/transformers/internal/generation_utils#logitsprocessor, allowing the model to avoid certain tokens by assigning their scores to minus infinity. Specifically, the model is restricted from generating n-grams absent from the source documents.
Additionally, to optimize the models for producing concise and informative highlights, we impose additional constraints, including setting minimum thresholds on the number of words in each highlight and the overall number of highlights generated. Based on hyperparameter-tuning we set a minimum of three words for each highlight in both MDS and LFQA, and a minimum number of 30 highlights for MDS and 5 highlights for LFQA. These constraints contribute significantly to optimizing the performance and reliability of the fine-tuned generation models.
Lastly, in the planning component, the model organizes and forms clusters, each encapsulating highlights from the given input. To prevent excessively large clusters, we impose a maximum threshold of two highlights per cluster. This facilitates the generation of coherent sentences from each cluster.
@@ Appendix D Full-sentence AutoAIS
[…]
Table D.1: MDS results comparing our reported AutoAIS to a variant of AutoAIS where the full-sentences containing the highlights are used as premise. The difference between the two AutoAIS variants is that the former might contain non-grammatical sentences.
[…]
Table D.2: QA results comparing our two AutoAIS variants, similar to Table D.1.
Tables D.2 and D.1 present the results on the full-sentence AutoAIS vs. regular AutoAIS
@@
Figure D.1: A screenshot from the first step in our human evaluation annotation interface for the LFQA setting, which is a replication of the interface presented by Liu et al. (2023). The annotator is presented with a query and a response and has to rate two aspects of the response on a Likert scale between 1 to 5. The first aspect is the fluency and cohesiveness of the response. The second is the helpfulness of the response.
@@
Figure D.2: A screenshot from the second step in our human evaluation annotation interface for the LFQA setting, which is an adaption of the interface suggested by Liu et al. (2023). The annotator is presented with citations provided by the model. For each sentence, they have to rate the quality of the citations. Highlights are presented if available, with arrows to navigate between different highlights. Completed sentences are presented with a green arrow to indicate that the annotator can continue to the next sentence.
Appendix E Human Study
@@
For our human study, we followed the Controlled Crowdsourcing Methodology Roit et al. (2020) to detect qualified annotators via Amazon Mechanical Turk.191919www.mturk.com We reached out to potential candidates, who had exhibited high proficiency in similar tasks in the past, through email invitations sent from the MTurk platform. During the onboarding process, candidates reviewed the annotation guidelines and completed a closed qualification round consisting of five annotation tasks. This process, overseen by an author of the paper who provided individualized feedback, required approximately 20 minutes to complete, with participants compensated 4.5\$ for their efforts. Ultimately, six out of eleven candidates were selected for the actual evaluations, with compensation rates set at 2\$ per ALCE instance and 1\$ otherwise, reflecting the varying complexity of the instances. To ensure fair compensation, we also monitored the actual time required to complete each instance and offered post-hoc reimbursements when the required time exceeded our initial estimates by more than 5%.
@@
Moreover, to evaluate the consistency of attribution annotations, we computed the inter-annotator agreement (IAA) using Fleiss’ Kappa coefficient (Fleiss, 1971), based on a sample of 101 sentences and their attributions from 23 instances across both models, each annotated by 3 workers. [^20] We treat the annotators’ scores as binary, as described in Appendix F. The resulting Kappa coefficient was 0.37, indicating a fair level of agreement (Landis and Koch, 1977).
Appendix F Annotation Interface
@@
Screenshots from our annotation interface are presented in Figs. D.1 and D.2. Our interface mirrors the design from Liu et al. (2023), adapted to support citations with highlights. In addition, the documents are presented within the interface rather than via external links. Within this interface, annotators assess each sentence to determine the level of support provided by its accompanying citations, choosing one of the following options: “Full Support”, “Partial Support”, “No Support”, “Citation Has Support but also Refutes Statement”, “Statement is Unclear, Can’t Make Judgement”. Finally, when calculating the final AIS score, we treat instances labeled as “Partial Support” as non-attributed, while instances classified as “Citation Has Support but also Refutes Statement” are labeled as attributed, akin to Gao et al. (2023a).
Appendix G Partially-attributing Examples
@@
Table G.1 demonstrates a few attribution examples labeled by human annotators as partially-supporting.
@@
Table G.1: Example predictions of the Attr. FirstCoT which were annotated with partial support. The information inferred from the words in bold is not supported by the highlights provided by the model, but is supported by sentences in close proximity.
@@
Table G.2: Example predictions of the ICL Attr. First model when prompted with prefix and without prefix.
Appendix H With Prefix Versus without Prefix
Table G.2 presents several instances from our test set when the iterative sentence-level fusion was only conditioned on each iteration’s cluster, versus when it was also conditioned on the prefix.

### v3

@@ C.1 Content Selection
In the ICL method, we present the model with the source texts and guide it to copy specific segments from those texts (see Figs. B.1 and B.5 for prompt examples in the MDS and LFQA settings, respectively). We use a designated “<SPAN_DELIM>” markup as a delimiter between such consecutive spans, for better parsing. Within the MDS framework, the model is tasked with identifying salient excerpts, whereas in the context of LFQA, the model is also provided with a query, and is asked to find the parts of the texts that are relevant to answering this question. Further, following our analyses of each setting’s development sets, we restrict the model to an average and maximum number of words to choose (an average of 200 and 100 words and a maximum of 900 and 200 words for MDS and LFQA, respectively). Lastly, to manage source texts that exceed the model’s input size limit, we truncate the fewest necessary sentences from those texts to ensure they fit within the specified limit. [^17] We use spaCy Honnibal and Montani (2017) to separate documents into sentences. After performing hyperparameter tuning, in the MDS we set $T=0.1$ and the number of in-context examples to 2, while in LFQA we set $T=0.3$ and the number of examples to 4.
@@
Examples are provided in Figs. B.9 and B.12. The input to the model is the set of all documents, concatenated with a special markup “<doc-sep>”, which was also used in pre-training in Primera. To fit the input in a context length of 4096, we trim each document to the length of 4096 divided by the number of documents in each example. The output is a set of highlights separated with a new special markup “<highlights-separator>”. We run hyperparameter-tuning on each setting based on the development set. The hyperparameters used for MDS are warmup steps of 229, weight decay of 0.5, and a learning rate of 4e-4. Conversely, the hyperparameters used for LFQA are warmup steps of 234, weight decay of 0, and a learning rate of 1e-4. We chose the best model checkpoint by calculating Intersection-over-Union (IoU) between the gold highlights spans and the predicted highlights spans.
@@ C.2 Generation Planning
In the ICL setting, we present the model with the source text where we insert the mark-ups “<highlight_start>” and “<highlight_end>” before and after each pre-selected span, (see Figs. B.2 and B.6 for the MDS and LFQA settings, respectively). We also enumerate the highlights across all documents, and for the LFQA setting, we also add the query. The model is directed to cluster highlighted segments that are likely to co-occur in a sentence in the same order it would then merge them into an output. We discovered that providing the model with task-specific motivation enhances its performance; for instance, in MDS, the objective is to fuse highlights into a summary, whereas in LFQA, it aims to formulate a response to the query. To facilitate the parsing of the model’s response, we encourage it to generate its response as a list of single-value dictionaries, with “cluster” as the key and a list of indices corresponding to the highlights within the current cluster as the value. For hyperparameter tuning, we divide the gold “highlights” into pairs. Pairs that correspond to the same target output sentence are categorized as positive, whereas others are considered negative. Subsequently, we evaluate the F1 score to assess the alignment between the predicted and gold clusters. This involves calculating the True Positives, False Negatives, and False Positives for the highlights pairs, based on the predicted clusters. Following our analyses, we opt for 2 examples and $T=0.5$ in the MDS setting, and 1 example and $T=0.1$ in the LFQA setting.
@@
Examples are provided in Figs. B.10 and B.13. There are two variants for generation planning: generative clustering and joint. In the generative clustering approach, the input to the model is the set of all highlights, concatenated with a special markup “<highlight-separator>”, which is the same as the output from the content selection model. The output is then a set of clusters, where each cluster is separated by a special markup “<cluster-separator>”, and inside each cluster we have a set of highlights in the same format as the input. We run hyperparameter-tuning on each setting based on the development set. The hyperparameters used for MDS are warmup steps 0, weight decay 0, and learning rate 5e-5. The hyperparameters used for LFQA are warmup steps 298, weight decay 0, and learning rate 2e-4. At inference time, we used a beam size of 1. We use the same logits processor described in Section C.1 to enforce the model to generate valid highlights.
@@ C.3 Attributed Fused Output
For the ICL variant, we iteratively show the model only the current cluster’s highlights, in the same manner, we do so in the generation planning ICL experiments (see Section C.2). Additionally, we present the model with what it generated to that point and guide it to generate the next sentence in the paragraph, which should cover all and only the highlighted spans (see Figs. B.3 and B.7). We use 3 examples and $T=0.3$ for the MDS setting, and 2 examples with $T=0.1$ for the LFQA setting.
For the Chain-of-Thought approach, which merges generation planning with the fusion steps, we present the model with the “highlighted” source texts and ask it to fuse those highlights into a coherent output. Additionally, we guide the model, both in the instructions and through in-context demonstrations to iteratively cluster highlights and fuse them into a sentence, before generating the final output (see Figs. B.4 and B.8). We find that 3 in-context examples and $T=0.1$ work best for MDS, while for LFQA it is 1 example and $T=0.3$ that yield the best results on the development set.
@@
Examples are provided in Figs. B.11 and B.14. The input to the attributable fusion model is a single cluster of highlights, where highlights are separated by a special markup. The output is then a single sentence of the response. This model is called iteratively, and at each iteration, we include the prefix of the already generated response. The output also includes citations for every sentence in the generated response based on the highlights used to generate the particular sentence. The hyperparameters for the fusion model are for MDS are warmup steps 54, weight decay 0.2, and learning rate 1e-4. The hyperparameters used for LFQA are warmup steps 175, weight decay 0.5, and learning rate 2e-5.
@@ C.4 End-to-end Generation
We include as a baseline end-to-end generation without decomposing the process into several steps.
@@
For the ICL setting, we instruct the model to perform the corresponding setting’s task, namely summarization for MDS, and answering a query based on the supplied document for LFQA. In addition to the varying number of in-context examples, we also evaluate the model in a zero-shot setting. Our analyses show that for MDS, 1 example and $T=0.5$ give the best performance, while for LFQA 3 examples and $T=0.7$ are optimal.
@@ C.6 Constrained Decoding
Both the content selection component and the planning component are designed to exactly copy text spans from source documents. To guide the generation to precisely follow the lexical structure of the input’s text, we employ a constrained decoding strategy for the fine-tuned models. This strategy relies on a logits processor [^18] https://huggingface.co/docs/transformers/internal/generation_utils#logitsprocessor, allowing the model to avoid certain tokens by assigning their scores to minus infinity. Specifically, the model is restricted from generating n-grams absent from the source documents. Additionally, to optimize the models for producing concise and informative highlights, we impose additional constraints, including setting minimum thresholds on the number of words in each highlight and the overall number of highlights generated. Based on hyperparameter-tuning we set a minimum of three words for each highlight in both MDS and LFQA, and a minimum number of 30 highlights for MDS and 5 highlights for LFQA. These constraints contribute significantly to optimizing the performance and reliability of the fine-tuned generation models. Lastly, in the planning component, the model organizes and forms clusters, each encapsulating highlights from the given input. To prevent excessively large clusters, we impose a maximum threshold of two highlights per cluster. This facilitates the generation of coherent sentences from each cluster.
[…]
Table C.1: MDS results comparing our reported AutoAIS to a variant of AutoAIS where the full-sentences containing the highlights are used as premise. The difference between the two AutoAIS variants is that the former might contain non-grammatical sentences.
[…]
Table C.2: LFQA results comparing our reported AutoAIS to a variant of AutoAIS where the full-sentences containing the highlights are used as premise. The difference between the two AutoAIS variants is that the former might contain non-grammatical sentences.
@@ Appendix D Full-sentence AutoAIS
Tables C.2 and C.1 present the results on the full-sentence AutoAIS vs. regular AutoAIS
Prediction 1 Generated sentence: Hawaii’s tourism industry is worth \$14 billion, and hotels and resorts are offering special packages to attract same-sex couples. [1] [1] …There’s a large number of firms that are specialized in the marriage business." In Hawaii, tourism is a \$14 billion industry. Now, the island chain is positioning itself for a spike in visitors - among them, Honolulu-based hotel chain Aqua Hospitality. It already offers LGBT travel deals, including the "out and proud" package and the "civil unions romance special." Bigger beachfront resorts are also getting in on the action. The Sheraton Waikiki is the only hotel on Oahu providing on-site marriage licenses. General Manager Kelly Sanders says now the hotel is running a new ad campaign in mainland LGBT publications … 2 Generated sentence: Patten provided lobbying and consulting services to a Ukrainian oligarch and his political party, but failed to register as a foreign agent as required by law. [1,2] [1] …From 2014, Patten provided a "prominent" Ukrainian oligarch who isn’t named in court papers and his Opposition Bloc political party with lobbying and consulting services, according to the criminal information. A company Patten co-owned with a Russian national received more than \$1 million for the work, the U.S. said. Details about the Ukrainian oligarch match those of Sergei Lyovochkin, an Opposition Bloc leader whom prosecutors have previously identified as funding Manafort’s work in Ukraine. He couldn’t immediately been reached for comment. As part of his lobbying work, Patten violated the Foreign Agents Registration Act by not disclosing the work to the U.S., prosecutors said. No date has been set for his sentencing … [2] …Prosecutors say Patten, who formed a consulting company with a person identified only as "Foreigner A," worked to set up meetings with members of Congress and also drafted talking points for Capitol Hill meetings. The goal, prosecutors say, was to influence U.S. policy, but they say Patten never filed under the Foreign Agents Registration Act. The law is aimed at promoting transparency about lobbying efforts in the United States. … 3 Generated sentence: On Tuesday, a shooting at a Home Depot store in North Dallas left three people injured: two police officers, Crystal Almeida and Rogelio Santander, and a Home Depot loss-prevention officer, Scott Painter. [1,2,3] [1]The wounded officer is Crystal Almeida, 26. The third victim was identified as Scott Painter, a Home Depot loss-prevention officer. Almeida and Painter were still in critical condition on Wednesday, but they were making "remarkable recoveries," Dallas Police Chief U. Renee Hall said at a news conference Wednesday morning. … [2] …Along with Officer Santander, Officer Crystal Almeida and the Home Depot security guard, now identified as Scott Painter, all underwent surgery after the shooting. Officer Almeida and Mr. Painter both remain hospitalized. During the press conference Chief Hall said, "We are happy to report that Officer Crystal Almeida and our loss-prevention officer Scott Painter is making remarkable recovery. They are still in critical condition but we are optimistic about what we’re seeing with them right now." The man who police say shot the officers and the security guard - Armando Luis Juarez - woke up behind bars today. After the shooting, a manhunt, and late night police chase, Juarez was taken into custody and ultimately transferred to the Dallas County Jail very early Wednesday morning. Officers Santander and Almeida both joined the Dallas Police Department there years ago, assigned to the Northeast Division. [3] …Mayor Mike Rawlings was presiding over a city council meeting when he announced the death of Rogelio Santander, a member of the police force for three years. Santander, officer Crystal Almeida and a loss-prevention officer for Home Depot were shot Tuesday by a man identified by police as 29-year-old Armando Luis Juarez. The two officers and the store loss-prevention officer underwent surgery for their injuries after the shooting in the north of the city, Dallas Police Chief U. Renee Hall said late Tuesday. Almeida and the loss-prevention officer, who hasn’t been identified, were in critical condition Wednesday.
Table D.1: Attribution examples for several sentences generated by the Attr. FirstCoT variant for the MDS settings.
Appendix E Fine-grained Attribution Examples
Table D.1 shows fine-grained attributions examples, generated by Attr. FirstCoT in the MDS setting.
@@ Appendix E Fine-grained Attribution Examples
Figure E.1: A screenshot from the first step in our human evaluation annotation interface for the LFQA setting, which is a replication of the interface presented by Liu et al. (2023). The annotator is presented with a query and a response and has to rate two aspects of the response on a Likert scale between 1 to 5. The first aspect is the fluency and cohesiveness of the response. The second is the helpfulness of the response.
@@
Figure E.2: A screenshot from the second step in our human evaluation annotation interface for the LFQA setting, which is an adaption of the interface suggested by Liu et al. (2023). The annotator is presented with citations provided by the model. For each sentence, they have to rate the quality of the citations. Highlights are presented if available, with arrows to navigate between different highlights. Completed sentences are presented with a green arrow to indicate that the annotator can continue to the next sentence.
Appendix F Human Study
@@
For our human study, we followed the Controlled Crowdsourcing Methodology Roit et al. (2020) to detect qualified annotators via Amazon Mechanical Turk.191919www.mturk.com We reached out to potential candidates, who had exhibited high proficiency in similar tasks in the past, through email invitations sent from the MTurk platform. During onboarding, candidates reviewed the annotation guidelines and completed a closed qualification round consisting of five annotation tasks. This process, overseen by an author of the paper who provided individualized feedback, required approximately 20 minutes to complete, with participants compensated 4.5\$ for their efforts. Ultimately, six out of eleven candidates were selected for the actual evaluations, with compensation rates set at 2\$ per ALCE instance and 1\$ otherwise, reflecting the varying complexity of the instances. To ensure fair compensation, we also monitored the actual time required to complete each instance and offered post-hoc reimbursements when this time exceeded our initial estimates by more than 5%.
@@
Moreover, to evaluate the consistency of attribution annotations, we computed the inter-annotator agreement (IAA) using Fleiss’ Kappa coefficient (Fleiss, 1971), based on a sample of 101 sentences and their attributions from 23 instances across both models, each annotated by 3 workers. [^20] We treat annotators’ scores as binary, akin to Appendix G. The resulting Kappa coefficient was 0.37, indicating a fair level of agreement (Landis and Koch, 1977).
Appendix G Annotation Interface
@@
Screenshots from our annotation interface are presented in Figs. E.1 and E.2. Our interface mirrors the design from Liu et al. (2023), adapted to support citations with highlights. In addition, the documents are presented within the interface rather than via external links. Within this interface, annotators assess each sentence to determine the level of support provided by its accompanying citations, choosing one of the following options: “Full Support”, “Partial Support”, “No Support”, “Citation Has Support but also Refutes Statement”, “Statement is Unclear, Can’t Make Judgement”. Finally, when calculating the final AIS score, we treat instances labeled as “Partial Support” as non-attributed, while instances classified as “Citation Has Support but also Refutes Statement” are labeled as attributed, akin to Gao et al. (2023a).
Appendix H Partially-attributing Examples
@@
Table H.1 shows a few attribution examples labeled by human annotators as partially-supporting.
@@
Table H.1: Example predictions of the Attr. FirstCoT which were annotated with partial support. The information inferred from the words in bold is not supported by the highlights provided by the model, but is supported by sentences in close proximity.
@@
Table H.2: Example predictions of the ICL Attr. First model when prompted with prefix and without prefix.
Content Selection Generation Planning Sentence Fusion CoT End-to-end ALCE
Input Output Input Output Input Output Input Output Input Output Input Output
ICL MDS
6879
($\pm$ 1893)
551
($\pm$ 467)
6425
($\pm$ 1728)
11
($\pm$ 22)
17003
($\pm$ 13203)
133
($\pm$ 80)
10273
($\pm$ 1558)
478
($\pm$ 220)
6879
($\pm$ 1893)
177
($\pm$ 55)
7341
($\pm$ 1974)
193
($\pm$ 44)
LFQA
10727
($\pm$ 3206)
142
($\pm$ 161)
2536
($\pm$ 1651)
9
($\pm$ 13)
3734
($\pm$ 1784)
45
($\pm$ 28)
2962
($\pm$ 635)
127
($\pm$ 89)
10727
($\pm$ 3206)
91
($\pm$ 85)
8476
($\pm$ 3809)
95
($\pm$ 89)
FT MDS
1661
($\pm$ 116)
456
($\pm$ 30)
1041
($\pm$ 75)
456
($\pm$30)
348
($\pm$ 7)
29
($\pm$ 0.6)
1661
($\pm$ 116)
283
($\pm$ 9)
LFQA
2883
($\pm$162)
121
($\pm$11)
348
($\pm$36)
121
($\pm$11)
209
($\pm$18)
27
(+ -1)
2883
($\pm$ 162)
60
($\pm$ 5)
Table H.3: Average number of tokens (and standard deviation) for each of the steps in our pipeline, as well as our two baselines.
Content Selection Generation Planning Sentence Fusion CoT End-to-end ALCE
ICL MDS 33.1 5.2 5.6 17.1 7.0 9.2
LFQA 6.3 4.7 6.8 3.3 6.1 6.7
FT MDS 68.8 5.7 2.4 - 1.7 -
LFQA 10.6 3.4 1.6 - 1.2 -
Table H.4: Average time (in seconds) each of the steps in our pipeline takes per instance, as well as for our two baselines.
Appendix I With Prefix Versus without Prefix
Table H.2 presents several instances from our test set when the iterative sentence-level fusion was only conditioned on each iteration’s cluster, versus when it was also conditioned on the prefix.
Appendix J Latency and Total Number of Input and Output Tokens
In Table H.3 we present the average number of tokens processed in the input and also the average number of generated tokens for each of our pipeline’s subtasks, and also for our baselines. Additionally, in Table H.4 we present the average time in seconds it takes to run each of our pipeline’s subtasks, as well as our baselines.
Appendix K Post-generation Attribution
We also investigated a post-generation attribution variant, where first an unattributed output was generated, and then relevant attribution was retrieved. To that end, we first generated the task-specific output, using our baseline from Section 4.2. Then, to obtain attributions, we adapted the RARR algorithm, as outlined by Gao et al. (2023a), to suit our context. In contrast to its original application of retrieving relevant documents, our adaptation employs the grounding texts directly. The process involves generating queries from each generated sentence about the various facets it encompasses, identifying 4-sentence paragraphs within the source texts that correspond to each query through a query-document relevance model, and answering query based on these paragraphs and the output sentence. The comparisons of these responses determine the selection of attributing paragraphs; only those that either entail or contradict are retained as valid attributions. The results, as detailed in Table K.1, show that this method produced both longer and less accurate attributions compared to those derived from our approach.
Task RougeL BertScore AutoAIS Length No Att. (%)
MDS 19.0 86.1 67.8 137.7 0.7
LFQA 33.1 89.6 50.2 149.0 9.6
Table K.1: Results of the post-generation attribution approach, adapted from the RARR algorithm, on the Multi-document summarization (MDS) and the Long-form Question-answering (LFQA) tasks.